qwevery:

The price of a car has been reduced from $25,000 to $21,750. What is the percentage decrease of the price of the car? A. 32.5% B. 87% C. 13% D. 1.3%

ByrdmanjrLakers:

It is 87%

ByrdmanjrLakers:

25000 * 87% = 21750

TETSXPREME:

Welcome To QuestionCove What do you think

TETSXPREME:

you have to divide

TETSXPREME:

@byrdmanjrlakers wrote:
It is 87%
dont give direct answers

qwevery:

it was wrong : First, find the difference between the two car prices. 25,000- 21,750=3,250 Then, divide by the original listing price, and convert to a percent. 3,250/25,000 = 0.13 =13% The percentage decrease of the price of the car is 13%.

qwevery:

thats what it said-

ByrdmanjrLakers:

ouch, shoot my bad

qwevery:

it fine lol

ByrdmanjrLakers:

@tetsxpreme wrote:
@byrdmanjrlakers wrote:
It is 87%
dont give direct answers
aight then kid

qwevery:

okay ima close now